# version 0.75

### v4

@@ 3.2.1 Markov Chains
It consists of a set of states and a transition matrix describing the probability of moving from one state to another. The probabilities of moving between states are the weights on the edges of the graph. The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$. Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as below
@@
    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \ |\ \ P_{ii}^{t}>0\right\},$
    is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor”.
The ergodic theorem asserts that every irreducible and aperiodic Markov chain has a unique stationary distribution. Its proof is a beautiful application of spectral graph theory. Exercise 3.6 asks the reader for a guided proof of the ergodic theorem.
@@
In exercise 3.6 we prove that a distribution that is reversible for a Markov chain is also stationary for this Markov chain.
@@ 3.2.2 Markov Reward Processes
Markov Reward Processes (MRPs) add the concept of reward to Markov Chains. At every state, the agent receives a reward, which we assume for simplicity to be in the range $[0,1]$. In terms of algorithmic consequences, no generality is lost by this assumption, and this is left as an exercise.
The cumulative reward is a sum of discounted rewards over time, so that immediate rewards are worth more than those received in the future. This is consistent with the economics literature, where a fixed amount of money is worth more today than in the future, due to positive interest rates.
@@ A Markov Reward Process is a tuple $(S,P,R,\gamma)$, where:
When the discount factor is strictly less than one, $G_{t}$ is well defined. In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that a stationary distribution for the underlying Markov chain exists, denoted by $\boldsymbol{\pi}$.
@@
The value function tells how valuable a state is as measured by the sum of its future, discounted rewards. One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@
In matrix form we can write the Bellman equation as $\mathbf{v}=\mathbf{\hat{R}}+\gamma P\mathbf{v}$, which simplifies to $\mathbf{v}=(I-\gamma P)^{-1}\mathbf{\hat{R}}$. Given $P$ and $\mathbf{\hat{R}}$, this is a linear system of equations, and thus the value function can be solved using Gaussian elimination in cubic time in the number of states.
@@ 3.2.3 Markov Decision Processes
The final layer of abstraction is to allow actions, or controls in the language of control theory. This gives the full Markov Decision Process model as defined below.
@@ A Markov Decision Process is a tuple $(S,P,R,A,\gamma)$ such that:
The central solution concept of study in the theory of reinforcement learning is called a policy. A policy is a mapping from state to distribution over actions, i.e.
@@
The goal of reinforcement learning is to determine a sequence of actions to maximize the long term reward. The Markovian structure of the problem suggests that this optimal value is realized by a stationary policy, i.e. there exists a fixed map from states to actions that maximizes the expected rewards.
@@
Similar to the value function of a Markov Reward Process, we can define the value function for a Markov Decision Process as follows.
@@
A fundamental theorem of reinforcement learning is the following.
@@
There exists a deterministic optimal policy $\pi^{*}$, which maps every state to a single action. All optimal policies achieve the same optimal value $\mathbf{v}^{*}(s)$ as well as $Q^{*}(s,a)$, and this holds for every state and action.
@@ 3.3 The Bellman Equation
A fundamental property of Markov Decision Processes is that they satisfy the Bellman equation for dynamic programming. The importance of this equation is paramount in algorithm design: we describe two efficient methods for reinforcement learning based on this equation later in this chapter. In later chapters we also make use of the Bellman equation to derive the stochastic-optimal LQR controller.
The Bellman equation for any policy $\pi$ relates the value function for that policy to itself via one step of application for that policy. Mathematically, this amounts to the following:
@@
    $R^{\pi}$ is the induced expected reward vector when playing policy $\pi$, i.e. ${R^{\pi}}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(s)}[R_{sa}]$.
@@
The Bellman equation gains prominance when applied to the optimal policy of a Markov Decision Process. It is then called the Bellman optimality equation, and states that for every state and action, we have
@@
where $Q^{*}$ is the $Q$ function of the optimal policy, i.e.
@@
In subsequent chapters, we will also consider cases where the return of interest is either the reward accumulated in a finite number of time steps, i.e. $G=\sum_{t=1}^{T}R_{t}$ in the finite-time setting, or the infinite-time average reward case, i.e. $G=\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}R_{t}$.
@@ 3.3.1 Reinforcement Learning as Linear Programming
Notice that this is a linear program! The number of variables is $|S|$, and the number of constraints is $|S|\times|A|$, which can be solved in polynomial time in the number of states and actions. The feasibility of this linear program is not immediate, but will be established in the next section.
@@ 3.4 The Value Iteration Algorithm
In the previous section we established the Bellman equation and its consequence to a polynomial time solution for reinforcement learning in Markov Decision Processes. In this section we study a fundamental algorithm for solving Markov Decision Processes that is a more direct application of dynamic programming and the Bellman equation.
@@
Define the residual vector $\mathbf{r}_{t}(s)=\mathbf{v}^{*}(s)-\mathbf{v}_{t}(s)$. Then the optimal value vector $\mathbf{v}^{*}$ is unique and value iteration converges to the optimal value function and policy at a geometric rate of
@@
Or in vector form, $\mathbf{r}_{t+1}\leq\gamma P^{\mathbf{a}^{*}}\mathbf{r}_{t}$. Furthermore, $\mathbf{r}_{t}\geq 0$ is entry-wise non-negative for all $t$. (See Exercise 3.6.) Therefore, it holds that
@@ 3.5 Bibliographic Remarks
The section on reinforcement learning in the text [RN02] introduces basic concepts, along with commentaries on related topics, such as the origins of the scalar reward model from utility theory. The classical text on reinforcement learning, authored by pioneers of the subject, is [SB18]. In addition, this text extensively discusses the early ventures into the field from researchers working in artificial intelligence, psychology, and neuroscience.
Another great reference is the lecture notes of [Sil15]. See [AJKS19] for an excellent introduction to the theoretical aspects of reinforcement learning, such as sample complexity bounds and reductions to supervised learning. The reader may also look to [Ber19] for an extensive survey of computational approaches to learning in unknown Markov decision processes.
@@
Although tabular MDPs fail to compactly represent stochastic linear dynamical systems, function-approximation schemes [JYWJ20, DKL+21, JKA+17, KKL+20] successfully capture the stochastic LQR setting.
@@ 3.6 Exercises
3.1. In this exercise we give a guided proof of the ergodic theorem. Consider a given irreducible and aperiodic Markov chain with transition matrix $P$.
@@
The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue equal to one is unique. We are going to assume this fact, and denote by $\mathbf{v}_{1}$ the corresponding eigenvector.
@@
3.5. Prove that the successive iterates $\mathbf{v}_{t}$ produced by the Value Iteration algorithm are monotonically increasing, i.e. $\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$, when initialized with $\mathbf{v}_{0}=\mathbf{0}$. Hence, conclude that for any $t$, $\mathbf{v}^{*}\geq\mathbf{v}_{t}$.
3.6. Let $P$ be an aperiodic and irreducible Markov chain. Prove that $P$ is reversible if and only if for any sequence of states $s_{1},\dots s_{n}$, it holds that
@@
Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories. Source: Wikipedia.
@@ 4.1 Approximating General Dynamics via LTV Systems
Consider any dynamics that can be expressed by a smooth function $f$. Such a system can be approximated by a time-varying linear system to an arbitrary accuracy determined by the time step interval between samples.
@@
where $J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is the Jacobian of $f$ with respect to the variables $\bar{\mathbf{x}}$. Writing the matrices ${A}_{t},{B}_{t}$ for the Jacobians, we have
@@
This derivation explains why in practice, the class of time varying linear dynamical systems is very expressive, at least for smooth dynamical systems. This is particularly true with hidden states that can capture complicated dynamics while imposing enough structure for efficient algorithms.
As we see next, stabilizability and controllability can be checked efficiently for linear systems, in contrast to general dynamics.
@@ 4.2 Stabilizability of Linear Systems
Recall that a system is stabilizable if from every starting state, there is a sequence of controls that drives the system to zero. Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows. Henceforth we denote by $\rho({A})$ the spectral radius of a matrix ${A}$, that is,
@@
the maximum modulus from all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can be characterized as follows.
@@
We henceforth prove sufficiency of this condition, and leave necessity as exercise 4.6. Both directions rely on the characterization of the spectral norm as follows:
@@
Suppose there exists a ${K}$ such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
@@ 4.2.1 Efficient Verification of Stabilizability
The spectral radius of a matrix can be computed to an arbitrary accuracy in polynomial time, and can be approximated by the power method. This implies that stability of a given system can be efficiently verified.
To efficiently check if a given system is stabilizable, we make note of the following result due to Lyapunov.
@@
Say there exists a matrix $P$ such that $P\succeq I$ and $P\succ A^{\top}PA$, then to show that $\rho(A)<1$, it suffices to prove that $\lim_{t\to\infty}A^{t}=0$. To observe the latter, first note that there exists a positive constant $c<1$ such that $cP\succeq A^{\top}PA$, implying that $c^{t}P\succeq(A^{t})^{\top}PA^{t}\succeq(A^{t})^{\top}A^{t}$ via repeated applications of the previous inequality. Therefore, we have that $\lim_{t\to\infty}(A^{t})^{\top}A^{t}=0$, and consequently that $\lim_{t\to\infty}A^{t}=0$ since $\|A^{\top}\|\|A\|=\|A^{\top}A\|$.
For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$. Since $P=I+A^{\top}PA$, $P$ satisfies the required conclusions. The fact that $P$ as defined exists, i.e. that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$. ∎
Using this observation, one can write a semi-definite mathematical program that checks if a given system is stabilizable, and if so produces a matrix $K$ that guarantees a spectral radius smaller than one.
@@
Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and it holds that $\rho((A+BK)^{\top})=\rho(A+BL)$. Now using this fact and Theorem 4.3, we know that stabilizability of a linear system is equivalent to the existence of a matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$. The above lemma allows us to reformulate this cubic constraint to a quadratic constraint. The stabilizability of a system is equivalent to existence of matrices ${K}$ and ${P}$ such that
@@
Lastly, we reparametrize this constraint as a linear semi-definite constraint as demonstrated below, by choosing ${L}={KP}$.
@@
4.3 Controllability of LDS
Recall that a system is controllable if and only if in the absence of noise and for every target state, there is a sequence of controls that drives the system state to that target. Linear dynamical systems exhibit a particularly elegant characterization of controllability. First, we need to define an object of central importance in linear dynamical system, the Kalman matrix.

### v5

@@ 3.2.1 Markov Chains
It consists of a set of states and a transition matrix that describes the probability of moving from one state to another. The probabilities of moving between states are the weights on the edges of the graph. The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$. Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as below
@@
    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \ |\ \ (P^{t})_{ii}>0\right\},$
    is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor".
The ergodic theorem states that every irreducible and aperiodic Markov chain has a unique stationary distribution. Its proof is a beautiful application of spectral graph theory. Exercise 3.6 asks the reader for a guided proof of the ergodic theorem.
@@
In exercise 3.6, we prove that a distribution that is reversible for a Markov chain is also stationary for this Markov chain.
@@ 3.2.2 Markov Reward Processes
Markov Reward Processes (MRPs) add the concept of reward to Markov Chains. In each state, the agent receives a reward, which, for simplicity, we assume to be in the range $[0,1]$. In terms of algorithmic consequences, no generality is lost due to this assumption, and this is left as an exercise.
The cumulative reward is the sum of discounted rewards over time, so that the immediate rewards are worth more than those received in the future. This is consistent with the economics literature, where a fixed amount of money is worth more today than in the future, because of positive interest rates.
@@ A Markov Reward Process is a tuple $(S,P,R,\gamma)$, where:
When the discount factor is strictly less than one, $G_{t}$ is well defined. In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that there is a stationary distribution for the underlying Markov chain, denoted by $\boldsymbol{\pi}$.
@@
The value function tells how valuable a state is as measured by the sum of its future discounted rewards. One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@
In matrix form, we can write the Bellman equation as $\mathbf{v}=\mathbf{\hat{R}}+\gamma P\mathbf{v}$, which simplifies to $\mathbf{v}=(I-\gamma P)^{-1}\mathbf{\hat{R}}$. Given $P$ and $\mathbf{\hat{R}}$, this is a linear system of equations, and thus the value function can be solved using Gaussian elimination in cubic time in the number of states.
@@ 3.2.3 Markov Decision Processes
The final layer of abstraction is to allow actions or controls in the language of control theory. This gives the full Markov Decision Process model as defined below.
@@ A Markov Decision Process is a tuple $(S,P,R,A,\gamma)$ such that:
The central solution concept of study in the theory of reinforcement learning is called a policy. A policy is a mapping from state to distribution over actions, that is,
@@
The goal of reinforcement learning is to determine a sequence of actions to maximize long-term reward. The Markovian structure of the problem suggests that this optimal value is realized by a stationary policy, i.e. there exists a fixed map from states to actions that maximizes the expected rewards.
@@
Similarly to the value function of a Markov Reward Process, we can define the value function for a Markov Decision Process as follows.
@@
A fundamental theorem of reinforcement learning is as follows.
@@
There exists a deterministic optimal policy $\pi^{*}$, which maps every state to a single action. All optimal policies achieve the same optimal value $\mathbf{v}^{*}(s)$ as well as $Q^{*}(s,a)$, and this holds for all states and actions.
@@ 3.3 The Bellman Equation
A fundamental property of Markov Decision Processes is that they satisfy the Bellman equation for dynamic programming. The importance of this equation is paramount in algorithm design: we describe two efficient methods for reinforcement learning based on this equation later in this chapter. In later chapters, we also make use of the Bellman equation to derive the stochastic-optimal LQR controller.
The Bellman equation for any policy $\pi$ relates the value function of that policy to itself through one step of application for that policy. Mathematically, this amounts to the following:
@@
    $R^{\pi}$ is the expected reward vector induced when playing the policy $\pi$, that is, ${R^{\pi}}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(s)}[R_{sa}]$.
@@
The Bellman equation gains prominence when applied to the optimal policy of a Markov Decision Process. Then it is called the Bellman optimality equation, and states that for every state and action, we have
@@
where $Q^{*}$ is the $Q$ function of the optimal policy, i.e.,
@@
In subsequent chapters, we will also consider cases where the return of interest is either the reward accumulated in a finite number of time steps, that is, $G=\sum_{t=1}^{T}R_{t}$ in the finite-time setting, or the infinite-time average reward case, i.e. $G=\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}R_{t}$.
@@ 3.3.1 Reinforcement Learning as Linear Programming
Notice that this is a linear program! The number of variables is $|S|$, and the number of constraints is $|S|\times|A|$, which can be solved in polynomial time in the number of states and actions. The feasibility of this linear program is not immediate but will be established in the next section.
@@ 3.4 The Value Iteration Algorithm
In the previous section, we established the Bellman equation and its consequence to a polynomial-time solution for reinforcement learning in Markov Decision Processes. In this section, we study a fundamental algorithm for solving Markov Decision Processes that is a more direct application of dynamic programming and the Bellman equation.
@@
Define the residual vector $\mathbf{r}_{t}(s)=\mathbf{v}^{*}(s)-\mathbf{v}_{t}(s)$. Then the optimal value vector $\mathbf{v}^{*}$ is unique and the value iteration converges to the optimal value function and policy at a geometric rate of
@@
Or in vector form, $\mathbf{r}_{t+1}\leq\gamma P^{\mathbf{a}^{*}}\mathbf{r}_{t}$. Furthermore, $\mathbf{r}_{t}\geq 0$ is entry-wise nonnegative for all $t$. (See Exercise 3.6.) Therefore, it holds that
@@ 3.5 Bibliographic Remarks
The section on reinforcement learning in the text [RN02] introduces basic concepts, along with commentaries on related topics, such as the origins of the scalar reward model from utility theory. The classical text on reinforcement learning, authored by pioneers of the subject, is [SB18]. In addition, this text extensively discusses early ventures into the field of researchers working in artificial intelligence, psychology, and neuroscience.
Another great reference is the set of lecture notes [Sil15]. See [AJKS19] for an excellent introduction to the theoretical aspects of reinforcement learning, such as sample complexity bounds and reductions to supervised learning. The reader may also refer to [Ber19] for an extensive survey of computational approaches to learning in unknown Markov decision processes.
@@
Although tabular MDPs do not compactly represent stochastic linear dynamical systems, function-approximation schemes [JYWJ20, DKL+21, JKA+17, KKL+20] successfully capture the stochastic LQR setting.
@@ 3.6 Exercises
3.1. In this exercise, we give a guided proof of the ergodic theorem. Consider an irreducible and aperiodic Markov chain with a transition matrix $P$.
@@
The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue equal to one is unique. We are going to assume this fact and denote by $\mathbf{v}_{1}$ the corresponding eigenvector.
@@
3.5. Prove that the successive iterates $\mathbf{v}_{t}$ produced by the Value Iteration algorithm increase monotonically, that is, $\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$, when initialized with $\mathbf{v}_{0}=\mathbf{0}$. Hence, conclude that for any $t$, $\mathbf{v}^{*}\geq\mathbf{v}_{t}$.
3.6. Let $P$ be an irreducible and aperiodic Markov chain. Prove that $P$ is reversible if and only if for any sequence of states $s_{1},\dots s_{n}$, it holds that
@@
Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories. Created by XaosBits using the Mathematica packages CurvesGraphics.nb by Gianluca Gorni and DrawGraphics.nb by David and Alice Park, licensed under CC-BY-2.5.
@@ 4.1 Approximating General Dynamics via LTV Systems
Consider any dynamics that can be expressed by a smooth function $f$. Such a system can be approximated by a time-varying linear system to an arbitrary accuracy determined by the time-step interval between samples.
@@
where $J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is the Jacobian of $f$ with respect to variables $\bar{\mathbf{x}}$. Writing the matrices ${A}_{t},{B}_{t}$ for the Jacobians, we have
@@
This derivation explains why in practice the class of time-varying linear dynamical systems is very expressive, at least for smooth dynamical systems. This is particularly true with hidden states that can capture complicated dynamics while imposing enough structure for efficient algorithms.
As we shall see next, stabilizability and controllability can be checked efficiently for linear systems, in contrast to general dynamics.
@@ 4.2 Stabilizability of Linear Systems
Recall that a system is stabilizable if, from every starting state, there is a sequence of controls that drives the system to zero. Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows. Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix ${A}$, that is,
@@
the maximum modulus of all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can be characterized as follows.
@@
We therefore prove the sufficiency of this condition and leave the necessity as exercise 4.6. Both directions rely on the characterization of the spectral norm as follows:
@@
Suppose that there exists a ${K}$ such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
@@ 4.2.1 Efficient Verification of Stabilizability
The spectral radius of a matrix can be computed to an arbitrary accuracy in polynomial time and can be approximated by the power method. This implies that the stability of a given system can be efficiently verified.
To efficiently check if a given system is stabilizable, we make note of the following result due to Lyapunov. We use the notation $A\succeq 0$ to denote that the matrix $A$ is positive semidefinite, meaning that it is symmetric and all its eigenvalues are nonnegative.
@@
Say there exists a matrix $P$ such that $P\succeq I$ and $P\succ A^{\top}PA$, then to show that $\rho(A)<1$, it suffices to prove that $\lim_{t\to\infty}A^{t}=0$. To observe the latter, first note that there exists a positive constant $c<1$ such that $cP\succeq A^{\top}PA$, implying that $c^{t}P\succeq(A^{t})^{\top}PA^{t}\succeq(A^{t})^{\top}A^{t}$ via repeated applications of the previous inequality. Therefore, we have $\lim_{t\to\infty}(A^{t})^{\top}A^{t}=0$, and consequently $\lim_{t\to\infty}A^{t}=0$ since $\|A^{\top}\|\|A\|=\|A^{\top}A\|$.
For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$. Since $P=I+A^{\top}PA$, $P$ satisfies the required conclusions. The fact that $P$ as defined exists, that is, that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$. ∎
Using this observation, one can write a semidefinite mathematical program that checks if a given system is stabilizable, and if so produces a matrix $K$ that guarantees a spectral radius smaller than one.
@@
Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and we have $\rho((A+BK)^{\top})=\rho(A+BK)$. Now using this fact and Theorem 4.3, we know that the stabilizability of a linear system is equivalent to the existence of matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$. The above lemma allows us to reformulate this cubic constraint to a quadratic constraint. The stabilizability of a system is equivalent to existence of matrices ${K}$ and ${P}$ such that
@@
Lastly, we reparameterize this constraint as a linear semidefinite constraint as demonstrated below, by choosing ${L}={KP}$.
@@
4.3 Controllability of Linear Dynamical Systems
Recall that a system is controllable if and only if in the absence of noise and for every target state there is a sequence of controls that drives the system state to that target. Linear dynamical systems exhibit a particularly elegant characterization of controllability. First, we need to define an object of central importance for linear dynamical systems, the Kalman matrix.